Need helps on parts B, C, E,and F

Costs of installing rooftop solar PV :

Calculate out-of-pocket costs and payback time for a rooftop solar installation. Assume that this PV system measures 6 m x 5 m and that the panels are 20% efficient.

(a) Peak power rating: The solar industry describes the “peak wattage” of a PV system as its power output at a hypothetical, industry-standard irradiance of 1000 W/m2. What would be the peak wattage (in W) of this PV system?

Answer:6000W

(b) Total installation costs: Peak wattage is used by the industry to determine installation costs. Assume that total installation costs (including the panels, parts, permits, and labor) are $2.80/W where the wattage refers to peak wattage. How much will this PV system cost to install?

(c) Out-of-pocket costs: Fortunately, the government has programs to reduce this cost. The federal tax credit for solar PV is 30% of installation (in other words, the U.S. government reimburses homeowners for 30% of the above cost). In addition, Arizona provides an additional tax rebate of $1000. When these one-time tax credits are deducted from the installation costs, what are the net out-of-pocket costs for this PV system?

(d) Annual generation:   Phoenix’s average daily insolation is 5.38 kWh/m2/day. Assume that this energy is entirely captured by the panels (in other words, don’t worry about the pitch of the roof or the direction of the panels). What is the annual electricity output (in kWh) of this PV system?

(e) Payback time: If Arizona’s electricity price is 12 cents/kWh, what is the payback time of this PV system (in years to two decimal places)? Hint: think about the electricity generated by the PV system as “saving” you the cost of buying electricity from the grid.

(f) Geographic variation: Now compare to another geographic location: the great city of Portland, OR. Electricity there is cheaper at 10 cents/kWh. The state of Oregon currently offers a rebate of $1.30/W of installed capacity (where installed capacity refers to the installation’s peak wattage) up to a $6,000 maximum per system. Assume the same panel efficiency, installation costs, and federal tax credit as we used for Arizona. The Portland, OR's annual average daily insolation is 3.82 kWh/m2/day. Based on these parameters, in Portland, OR what would be this system’s:

i. Out-of-pocket installation cost (in $)?

ii. Annual electricity generation (in kWh)?

iii. Payback time (in years)?
Answer #1
(b) Total installation costs
A Peak wattage 6000 Watt
B Installation cost per watt $2.80
C=A*B Total installation costs $16,800
.(c) Out-of-pocket costs
D=0.3C Tax credit $5,040
E additional tax rebate by Arizona $1,000
F=C-D-E Out-of-pocket costs $10,760
(d) Annual generation:
G Average daily insolation 5.38 KWh/sqm/day
H Efficiency 20%
I Area in sq metre 30 (6*5)
J=G*H*I Generation per day 32.28 KWh
K=J*365 Annual generation: 11782.2 KWh
(e) Payback time
L Cost of electricity per KWh $0.12
M=K*L Annual savings $1,413.86
N=F/M Payback time in years 7.61
(f) Geographic variation
(i) Out-of-pocket costs
Tax credit $5,040
additional tax rebate by Portland $6,000
Out-of-pocket costs $5,760 (16800-5040-6000)
(ii) Annual generation:
Average daily insolation 3.82 KWh/sqm/day
Efficiency 20%
Area in sq metre 30
Generation per day 22.92 KWh (3.82*30*0.2)
Annual generation: 8365.8 KWh (22.92*365)
(iii) Payback time
Cost of electricity per KWh $0.10
Annual savings $836.58 (0.10*8365.8)
Payback time in years 6.89 (5760/836.58)